# Entropic Confinement and Mode Connectivity in Overparameterized Neural Networks

### v1

@@ A.1 $k$-step projected SGD
In order to directly measure the effect of entropic forces in a controlled setting, we use a modified version of SGD. Our algorithm deals with two conflicting considerations: First, we would like to limit the scope of our observation to models that lie on a linear path, or more generally models that lie on a MEP. However, we would also like to run optimization in such a way that entropic forces arising from curvature are still relevant to the optimization dynamics. The key observation is that if we were to run SGD on a line in parameter space, projecting back to the line after each optimization step, we would remove the effect of entropic forces, which arise from noisy multi-step optimization dynamics wei2019noiseaffectshessianspectrum. Motivated by this, we propose a natural algorithm that trades off between these two considerations by taking multiple SGD steps between before projecting the parameters back to the liner path (or MEP).
@@ A.3 Supplementary Figures
Figure 7: Properties of MEPs constructed with AutoNEB. The AutoNEB algorithm updates the positions of pivots without changing the lengths $\ell$ of the segments connecting them, and adds new pivots along segments where the loss is not well-approximated by the linear interpolation, see pmlr-v80-draxler18a. As a result, the Euclidean distance between consecutive pivots is not constant, and pivots tend to be denser near the middle of the MEP. (A) Relative Euclidean distance of each pivot measured from the first pivot, illustrating the cumulative distance along the MEP. (B) Lengths of individual segments between consecutive pivots, showing that the spacing $\Delta\ell$ is non-uniform. Note that the values in (A) correspond to the cumulative sum of the segment lengths shown in (B).

### v2

@@ A.1 $k$-step projected SGD
In order to directly measure the effect of entropic forces in a controlled setting, we use a modified version of SGD. Our algorithm deals with two conflicting considerations: First, we would like to limit the scope of our observation to models that lie on a linear path, or more generally models that lie on a MEP. However, we would also like to run optimization in such a way that entropic forces arising from curvature are still relevant to the optimization dynamics. The key observation is that if we were to run SGD on a line in parameter space, projecting back to the line after each optimization step, we would remove the effect of entropic forces, which arise from noisy multi-step optimization dynamics Wei and Schwab (2019). Motivated by this, we propose a natural algorithm that trades off between these two considerations by taking multiple SGD steps between before projecting the parameters back to the liner path (or MEP).
@@ A.3 Supplementary Figures
Figure 7: Properties of MEPs constructed with AutoNEB. The AutoNEB algorithm updates the positions of pivots without changing the lengths $\ell$ of the segments connecting them, and adds new pivots along segments where the loss is not well-approximated by the linear interpolation, see Draxler et al. (2018). As a result, the Euclidean distance between consecutive pivots is not constant, and pivots tend to be denser near the middle of the MEP. (A) Relative Euclidean distance of each pivot measured from the first pivot, illustrating the cumulative distance along the MEP. (B) Lengths of individual segments between consecutive pivots, showing that the spacing $\Delta\ell$ is non-uniform. Note that the values in (A) correspond to the cumulative sum of the segment lengths shown in (B).